Rank the following three single taxpayers in order of the magnitude of taxable income (from lowest to highest). (First mean hThrough November, Cameron has received gross income of $52,500. For December, Cameron is considering whether to accept one moPlease Bold answers, thank you!
Answer #1

Amounts are in $

Part 1 :

Taxable income = Gross Income - Deduction for AGI - Itemized deduction - QBI deduction

1. Ahmed = 96,100 - 15,550 = 80,550

2. Baker = 96,100 - 8,550 - 7,650 - 1,925 = 77,975

3. Chin = 96,100 - 14,325 - 10,000 = 71,775

Part 2 :

Assuming for this question, the year is 2019

Description Assumption 1 Assumption 2
Gross income before new work engagement 52,500 52,500
Income from engagement 7,450 6,750
Additional for AGI deduction (3,550) 0
Adjusted gross income 56,400 59,250
Standard deduction (12,200) (12,200)
Deduction for QBI 0 (1,350)
Taxable Income 44,200 45,700

Note :

Standard deduction for single tax payer in year 2019 is $12,200

(If you assume any other year, you will have different Standard deduction amount and this different taxable income)

(Given there are no itemized deduction available)

QBI deduction is 20% of QBI income = 20% x 6,750 = 1,350